OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (johnweldon1993):

set up the point slope again y - (-5) = -3(x-2) y+5 = -3x + 6 y = -3x + 1 :) hope I'm helping :)

OpenStudy (anonymous):

You are! Thank you. (: